# A perturbative construction of primitive forms from log Landau-Ginzburg mirrors of toric manifolds

### v1

@@ 3.3. Logarithmic Frobenius manifold
Since $\varphi_{1},\cdots,\varphi_{\mu}$ can also be viewed as a basis of sections of $\mathcal{H}_{F}$, there exists an ^$\hat{\mathscr{R}}((u))$-valued matrix $A$ such that
@@
Let $\mathcal{H}_{F,-}$ be the ^$\hat{\mathscr{R}}[u^{-1}]$-submodule of $\mathcal{H}_{F,\pm}$ generated by $u^{-1}DE(\varphi_{1})$, normal-$\cdots$, $u^{-1}DE(\varphi_{\mu})$. Then $\mathcal{H}_{F,-}$ defines a good opposite filtration, in other words, we have
@@
To construct a primitive form, we solve the equation $A\equiv CB^{-1}\ (\text{mod }\mathcal{I}^{k+1})$ by induction on $k$, so that we can write $A=CB^{-1}$, where $C$ is an ^$\hat{\mathscr{R}}[[u]]$-valued matrix and $B$ is of the form $\text{Id}_{\mu\times\mu}+B^{\prime}$, in which $B^{\prime}$ is an $u^{-1}\hat{\mathscr{R}}[u^{-1}]$-valued matrix. Hence
@@
Denote $(\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{\mu}):=(DE(\varphi_{1}),\cdots,DE(\varphi_{\mu}))B$. Then $-^\operatorname{K}_{F}(\tilde{\varphi}_{i},\tilde{\varphi}_{j})\in\hat{\mathscr{R}}$.
@@
Since $C$ is ^$\hat{\mathscr{R}}[[u]]$-valued and $\operatorname{K}_{F}(\varphi_{i},\varphi_{j})\in\hat{\mathscr{R}}[[u]]$, we see that $\operatorname{K}_{F}(\tilde{\varphi}_{i},\tilde{\varphi}_{j})\in\hat{\mathscr{R}}[[u]]$. On the other hand, by Lemma 3.8 and the form of $B$, we have $\operatorname{K}_{F}(\tilde{\varphi}_{i},\tilde{\varphi}_{j})\in\hat{\mathscr{R}}[u^{-1}]$. The result follows. ∎
@@ 3.4. Log LG mirrors of semi-Fano toric manifolds
To begin with, we arrange the good basis $\varphi_{1},\dots,\varphi_{\mu}$ (in Theorem 2.21) so that $\varphi_{1}=1$ (i.e. of weight degree 00), $\varphi_{2},\dots,\varphi_{\nu}$ are of weight degree $2$, while the remaining $\varphi_{j}$’s are of higher weight degrees. Let $\mathbf{F}$ be the vector space spanned by the basis
@@
[…]
@@
[…]
@@
[…]
where $\zeta$ is the primitive form as in Proposition 3.12.
@@ 3.4.1. The period map for log LG mirror of a semi-Fano toric manifold
We will study the period map restricted to $\hat{S}$ for the family $(X^{{\dagger}},F)$ mirror to a semi-Fano toric manifold (resp. Fano toric manifold) $Y$, which refers to those with all $p\in P$ having non-negative (resp. positive) weight degrees. We denote by $P_{0}$ the submonoid consisting of those $p$’s with weight degree 00.
@@
We take the basis $\{\varphi_{1},\dots,\varphi_{\mu}\}$ for $\mathcal{H}_{f,+}$ as in the beginning of this subsection, lift it to a frame of $\mathcal{H}_{F,+}$ using the same monoial expressions and call them $\psi_{i}$’s.
@@
Let $(X^{{\dagger}},F)$ be the log LG mirror of a semi-Fano toric manifold $Y$. Then the connection matrix $u^{-1}A$, as a matrix of $1$-forms in $\Omega_{S^{{\dagger}}}^{1}[[u]]$ written as a block matrix with respect to the frames $\psi_{1}$ in the first block, $\psi_{2},\dots,\psi_{\nu}$ in the second block and $\psi_{\nu+1},\dots,\psi_{\mu}$ in the third block, is of the form
@@
For $\psi_{1}=1$, we have $\nabla_{n}1=\sum_{s=1}^{d}u^{-1}z_{s}\iota_{n}(d\log z_{s})$ and we can express each $z_{s}$ in terms of $\psi_{2},\dots,\psi_{\nu}$ using relations in $\mathcal{H}_{F,\pm}$ as in the Moving Lemma 2.20. For other $\psi_{i}=P(\tau_{i})=z_{i_{1}}\cdots z_{i_{k}}$ which is identified with $P(\tau_{i})\Omega\in\mathcal{H}^{F}_{+}$, the action of the connection $\nabla_{n}$ is given by
@@
[…]
@@
where each $g_{i(-1)}(z^{p})$ is a formal series of $z^{p}$ where $p$’s have weight degree 00.
@@
Recall that in equation (15), we have $g_{ik}=0=h_{jk}$ for $k\geq 0$, $f_{k}=0$ for $k>0$ and $h_{j(-1)}=0$. The only possibility that cannot be ruled out by a degree argument is when $f_{0}$ and $g_{i(-1)}$ are series in $z^{p}$ where $p$’s have weight degree 00. The presence of $g_{i(-1)}$ contributes to the corresponding terms in $\log(\tau_{i})$, so we simply have to rule out the presence of $f_{0}$ and $f_{-1}$. Once again we have to look at the expression of $\psi_{1}$ in terms of $DE(\varphi_{i})$’s. Fix a decreasing sequence of ideals $\mathcal{J}_{i}$ containing $\mathcal{I}_{>2}$ such that $\mathcal{J}_{1}=m$, $\mathcal{J}_{i}/\mathcal{J}_{i+1}$ is one-dimensional and generated by $z^{p}$ for some $p$, and their intersection is $\mathcal{I}_{>2}$. Such a sequence can be found because $P$ is a toric monoid, using arguments similar to those in [8, below Lemma 4.17]. We define a $\mu\times\mu$ matrix $G$ by the change of frames $[\psi_{1},\dots,\psi_{\mu}]=[DE(\varphi_{1}),\dots,DE(\varphi_{\mu})]G$. Passing to the quotient $R((u))/\mathcal{J}_{i}((u))$, we write the corresponding matrix as $G_{i}$, which are compatible for different $i$’s. Notice that $G_{1}=\text{id}$. Write $\Omega_{\hat{S}^{{\dagger}}}^{1}\cong\hat{R}\otimes_{\mathbb{C}}V$ with elements in $V$ written as $d\log q_{i}$’s. The connection matrix when restricting to $H_{f,\pm}$ with respect to the basis $\varphi_{i}$’s is written as $u^{-1}N$, where $N$ is consists of constant $1$-forms with coefficient in $V$. Treating $N$ as a matrix of $1$-forms in $\Omega_{\hat{S}^{{\dagger}}}^{1}$, it gives the connection $1$-form with respect to the frame $DE(\varphi_{i})$’s by our construction. We have the change of frames relation
@@
when written as a block matrix as in Lemma 3.15. We can write $G_{i}$ as a matrix with coefficients in $R((u))\setminus\mathcal{J}_{i}((u))$, and relate $G_{i+1}=G_{i}+z^{p}\tilde{G}$ for $z^{p}\in\mathcal{J}_{i}\setminus\mathcal{J}_{i+1}$ and some $\tilde{G}$ with coefficients in $\mathbb{C}((u))$. Similarly we write $A_{i+1}=A_{i}+z^{p}\tilde{A}$ for the matrix $A$. Suppose the claim is true for $G_{i}$, and we consider the change of frames equation modulo $R((u))/\mathcal{J}_{i+1}$. We obtain
$(G_{i}+z^{p}\tilde{G})(A_{i}+z^{p}\tilde{A})=ud(G_{i}+z^{p}\tilde{G})+N(G_{i}+z^{p}\tilde{G}),$
@@ 3.4.2. Explicit computations for the Hirzebruch surface $\mathbb{F}_{2}$
Associated to $\mathbb{F}_{2}$ is the fan $\Sigma$ in $N_{\mathbb{R}}=\mathbb{R}^{2}$ whose $1$-dimensional cones are generated by $e_{1}=(1,0)$, $e_{2}=(0,1)$, $e_{3}=(-1,2)$ and $e_{4}=(0,-1)$. Let $D_{i}\subset\mathbb{F}_{2}$ be the toric divisor corresponding to the ray spanned by $e_{i}$. The monoid $P$ is isomorphic to $\mathbb{N}^{2}$ with generators $C_{1}$ and $C_{2}$ corresponding to rational curves lying in $D_{1}$ and $D_{2}$ respectively. Here $C_{1}$ has Chern number $2$ and $C_{2}$ has Chern number 00. The universal piecewise linear function $\varphi:|\Sigma|\rightarrow\mathbb{R}^{2}$ is defined by
@@
Letting $q_{1}=z^{(1,0)}$ and $q_{2}=z^{(0,1)}$, we see that $q_{1}$ has weight degree $2$ and $q_{2}$ has weight degree 00. $X^{\dagger}$ is given as a subvariety of $\text{Spec}(\mathbb{C}[q_{1},q_{2}][z_{1},z_{2},z_{3},z_{4}])$ by the relations
@@
To compare with known results, we will only compute the canonical coordinates and the primitive form when restricted to $\hat{R}=\mathbb{C}[[q_{1},q_{2}]]$ (the small quantum variables), ignoring the other two parameters $t_{1},t_{2}$. The Landau-Ginzburg superpotential is given by $F=z_{1}+z_{2}+z_{3}+z_{4}$. The sheaf $\Omega_{\hat{X}^{{\dagger}}/\hat{S}^{{\dagger}}}^{*}$ of log differential forms is locally free with generators $d\log z_{1}$, $d\log z_{2}$, and the holomorphic volume form is taken to be $\Omega=d\log z_{1}\wedge d\log z_{2}$. The cohomology $H^{*}(\Theta^{*}_{\hat{X}^{{\dagger}}/\hat{S}^{{\dagger}}}[[u]],u\partial+\{F,\cdot\})$ is concentrated at degree 00 which is generated by polynomials in $z_{i}$’s, $q_{i}$’s and $u$.
We have several useful relations such as:
where $f(q_{2})=\frac{1}{1-4q_{2}}$; the first two equalities come from the Moving Lemma 2.20, while the third equality is obtained from the following relations
@@
We arrange the maximal cones $\sigma_{i}$’s in $\Sigma$ in the counter-clockwise ordering so that $\sigma_{1}=\mathbb{R}_{\geq 0}e_{1}+\mathbb{R}_{\geq 0}e_{2}$. This gives us frames $1$, $z_{3}$, $z_{4}$ and $z_{1}z_{4}$ in $H_{F,+}$ which restrict to $\varphi_{i}$’s in $H_{f,+}$ after modulo $q_{1},q_{2}$, according to the choice made after equation (6).
@@
Next we solve for the flat sections $DE(\varphi_{i})$’s. We first restrict ourselves to $q_{1}=0$ and then extend to flat sections in $q_{1}$. Let $\psi_{i}$’s be the restrictions of $DE(\varphi_{i})$’s to $q_{1}=0$. Then we can simply take $\psi_{3}=z_{4}$ and $\psi_{4}=z_{1}z_{4}$. It remains to solve for $\psi_{1}$ and $\psi_{2}$. To do so, we let $\psi_{2}=a(q_{2})z_{3}+b(q_{2})z_{4}+c(q_{2})z_{1}z_{4}$. Then the equation $\nabla_{\frac{\partial}{\partial\log q_{2}}}\psi_{2}=0$ gives us
@@
with $\delta_{i}=\delta_{i}(q_{1},q_{2},u)=\sum_{j=-\infty}^{\infty}\delta_{ij}(q_{1},q_{2})u^{j}$ such that $\delta_{ij}(q_{1},q_{2})\in\mathbb{C}[[q_{1},q_{2}]]$ and the summation starts from $j=-N_{k}$ if we modulo $(q_{1},q_{2})^{k+1}$. The weight degrees of $DE(\varphi_{1})$, $DE(\varphi_{2})$, $DE(\varphi_{3})$ and $DE(\varphi_{4})$ are 00, $2$, $2$ and $4$ respectively. From these, we know that $\delta_{1}$ has weight degree 00, $\delta_{2}$ and $\delta_{3}$ both have weight degree $-2$ and $\delta_{4}$ has weight degree $-4$. Since the weight degree of $q_{1}$ is $4$ and that of $q_{2}$ is 00, we see that $\delta_{i,j\geq 0}=0$ for $i>0$, $\delta_{0,j>0}=0$ and $\delta_{0,0}$ is independent of $q_{1}$. Modulo $q_{1}$ we find that $\delta_{0,0}\cong 1$ from the expression of $\psi_{1}$. All in all, we see that the primitive form is simply given by $\zeta=1$.

### v3

@@ 3.3. Logarithmic Frobenius manifold
Since $\varphi_{1},\cdots,\varphi_{\mu}$ can also be viewed as a basis of sections of $\mathcal{H}_{F}$, there exists an $\hat{\mathscr{R}}((u))$-valued matrix $A$ such that
@@
Let $\mathcal{H}_{F,-}$ be the $\hat{\mathscr{R}}[u^{-1}]$-submodule of $\mathcal{H}_{F,\pm}$ generated by $u^{-1}DE(\varphi_{1})$, $\cdots$, $u^{-1}DE(\varphi_{\mu})$. Then $\mathcal{H}_{F,-}$ defines a good opposite filtration, in other words, we have
@@
To construct a primitive form, we consider a Birkhoff factorization of the $\hat{\mathscr{R}}((u))$-valued matrix $A$ in the proof of Lemma 3.7. Namely, we solve the equation $A\equiv CB^{-1}\ (\text{mod }\mathcal{I}^{k+1})$ by induction on $k$, so that $C$ is an $\hat{\mathscr{R}}[[u]]$-valued matrix and $B$ is of the form $\text{Id}_{\mu\times\mu}+B^{\prime}$, in which $B^{\prime}$ is an $u^{-1}\hat{\mathscr{R}}[u^{-1}]$-valued matrix (cf. [32, Sections 5.4 and 6]). Then we have
@@
Denote $(\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{\mu}):=(DE(\varphi_{1}),\cdots,DE(\varphi_{\mu}))B$. Then $\operatorname{K}_{F}(\tilde{\varphi}_{i},\tilde{\varphi}_{j})\in\hat{\mathscr{R}}$.
@@
Since $C$ is $\hat{\mathscr{R}}[[u]]$-valued and $\operatorname{K}_{F}(\varphi_{i},\varphi_{j})\in\hat{\mathscr{R}}[[u]]$, we see that $\operatorname{K}_{F}(\tilde{\varphi}_{i},\tilde{\varphi}_{j})\in\hat{\mathscr{R}}[[u]]$. On the other hand, by Lemma 3.8 and the form of $B$, we have $\operatorname{K}_{F}(\tilde{\varphi}_{i},\tilde{\varphi}_{j})\in\hat{\mathscr{R}}[u^{-1}]$. The result follows. ∎
@@ 3.4. Log LG mirrors of semi-Fano toric manifolds
To begin with, we arrange the good basis $\varphi_{1},\dots,\varphi_{\mu}$ (in Theorem 2.21) so that $\varphi_{1}=1$ (i.e. of weight degree 0), $\varphi_{2},\dots,\varphi_{\nu}$ are of weight degree $2$, while the remaining $\varphi_{j}$’s are of higher weight degrees. Let $\mathbf{F}$ be the vector space spanned by the basis
@@
We first describe the semi-infinite period map (or period map for short). We take the vector space $V=P^{gp}\times_{\mathbb{Z}}\mathbb{C}$ and treat an element $p\in P^{gp}\subset V$ as $\log z^{p}$. For $v\in\Theta_{\hat{S}^{{\dagger}}}\cong\hat{R}\otimes_{\mathbb{C}}V^{*}$, we declare its action on $\hat{R}_{V}=\hat{R}\hat{\otimes}_{\mathbb{C}}\text{Sym}^{*}V$ as a derivation using the natural pairing $\langle\cdot,\cdot\rangle$ between $V^{*}$ and $V$. Explicitly, if we write $\partial_{n}$ for $n\in V^{*}$, then we have $\partial_{n}(\log z^{p})=\langle n,p\rangle$ and $\partial_{n}(z^{p})=\langle n,p\rangle z^{p}$. Formally, we will extend the coefficient ring from $\mathscr{R}$ to $\mathscr{R}\hat{\otimes}_{\mathbb{C}}\text{Sym}^{*}V$ in order to write down the flat sections. We let $\mathbb{1}=DE(\varphi_{1})+\eta$ with $\eta\in\mathcal{H}_{F,\pm}\hat{\otimes}_{\mathbb{C}}\text{Sym}^{>0}V$ such that $\nabla_{v}\mathbb{1}=0$ $\forall v\in\Theta_{\hat{\mathscr{S}}}$.
@@
$\Psi:\hat{\mathscr{R}}\hat{\otimes}_{\mathbb{C}}\text{Sym}^{*}(V)\rightarrow\mathbb{C}[\tau_{1},\log\tau_{2},\dots,\log\tau_{\nu},\tau_{\nu+1},\dots,\tau_{\mu}]$
@@
$\Psi(z^{p},\log z^{p})=u[\zeta-\mathbb{1}]\in u\mathcal{H}_{F,-}/\mathcal{H}_{F,-}\cong\mathbf{F}\otimes_{\mathbb{C}}\hat{\mathscr{R}}\hat{\otimes}_{\mathbb{C}}\text{Sym}^{*}(V),$
where $\zeta$ is a primitive form as in Proposition 3.12.
@@ 3.4.1. The period map for log LG mirror of a semi-Fano toric manifold
We will study the period map restricted to $\hat{S}$ for the family $(X^{{\dagger}},F)$ mirror to a semi-Fano toric manifold (resp. Fano toric manifold) $Y$, which refers to those with all $p\in P$ having non-negative (resp. positive) weight degrees. We denote by $P_{0}$ the submonoid consisting of those $p$’s with weight degree 0.
@@
We take the basis $\{\varphi_{1},\dots,\varphi_{\mu}\}$ for $\mathcal{H}_{f,+}$ as in the beginning of this subsection, lift it to a frame of $\mathcal{H}_{F,+}$ using the same monomial expressions and call them $\psi_{i}$’s.
@@
Let $(X^{{\dagger}},F)$ be the log LG mirror of a semi-Fano toric manifold $Y$. Then the connection matrix $u^{-1}A$ of $\nabla_{n}$, as a $\Omega_{S^{{\dagger}}}^{1}[[u]]$-valued matrix and written as a block matrix with respect to the frames $\psi_{1}$ in the first block, $\psi_{2},\dots,\psi_{\nu}$ in the second block and $\psi_{\nu+1},\dots,\psi_{\mu}$ in the third block, is of the form
@@
For $\psi_{1}=1$, we have $\nabla_{n}1=\sum_{s=1}^{d}u^{-1}z_{s}\iota_{n}(d\log z_{s})$ and we can express each $z_{s}$ in terms of $\psi_{2},\dots,\psi_{\nu}$ using relations in $\mathcal{H}_{F,\pm}$ as in the Moving Lemma 2.20.
For other $\psi_{i}=P(\tau_{i})=z_{i_{1}}\cdots z_{i_{k}}$ which is identified with $P(\tau_{i})\Omega\in\mathcal{H}^{F}_{+}$, the action of the connection $\nabla_{n}$ is given by
@@
Let $(X^{{\dagger}},F)$ be the log LG mirror of a semi-Fano toric manifold. When restricted to the parameter space $\hat{R}\otimes_{\mathbb{C}}\text{Sym}^{*}V$, $\zeta=1$ is a primitive form and the period map takes the form
@@
where each $g_{i(-1)}(z^{p})$ is a formal series of $z^{p}$ where $p$’s have weight degree 0.
@@
Recall that in equation (16), we have $g_{ik}=0=h_{jk}$ for $k\geq 0$, $f_{k}=0$ for $k>0$ and $h_{j(-1)}=0$. The only possibility that cannot be ruled out by a degree argument is when $f_{0}$ and $g_{i(-1)}$ are series in $z^{p}$ where $p$’s have weight degree 0. The presence of $g_{i(-1)}$ contributes to the corresponding terms in $\log(\tau_{i})$, so we simply have to rule out the presence of $f_{0}$ and $f_{-1}$. Once again we have to look at the expression of $\psi_{1}$ in terms of $DE(\varphi_{i})$’s.
Fix a decreasing sequence of ideals $\mathcal{J}_{i}$ containing $\mathcal{I}_{>2}$ such that $\mathcal{J}_{1}=m$, $\mathcal{J}_{i}/\mathcal{J}_{i+1}$ is one-dimensional and generated by $z^{p}$ for some $p$, and their intersection is $\mathcal{I}_{>2}$. Such a sequence can be found because $P$ is a toric monoid, using arguments similar to those in [8, below Lemma 4.17]. We define a $\mu\times\mu$ matrix $G$ by the change of frames $\psi_{1},\dots,\psi_{\mu}]=[DE(\varphi_{1}),\dots,DE(\varphi_{\mu})]G$. Passing to the quotient $R((u))/\mathcal{J}_{i}((u))$, we write the corresponding matrix as $G_{i}$, which are compatible for different $i$’s. Notice that $G_{1}=\text{id}$.
Write $\Omega_{\hat{S}^{{\dagger}}}^{1}\cong\hat{R}\otimes_{\mathbb{C}}V$ with elements in $V$ written as $d\log q_{i}$’s. The connection matrix when restricting to $H_{f,\pm}$ with respect to the basis $\varphi_{i}$’s is written as $u^{-1}N$, where $N$ is consists of constant $1$-forms with coefficient in $V$. Treating $N$ as a matrix of $1$-forms in $\Omega_{\hat{S}^{{\dagger}}}^{1}$, it gives the connection $1$-form with respect to the frame $DE(\varphi_{i})$’s by our construction. We have the change of frames relation
@@
when written as a block matrix as in Lemma 3.15.
We can write $G_{i}$ as a matrix with coefficients in $R((u))\setminus\mathcal{J}_{i}((u))$, and relate $G_{i+1}=G_{i}+z^{p}\tilde{G}$ for $z^{p}\in\mathcal{J}_{i}\setminus\mathcal{J}_{i+1}$ and some $\tilde{G}$ with coefficients in $\mathbb{C}((u))$. Similarly we write $A_{i+1}=A_{i}+z^{p}\tilde{A}$ for the matrix $A$. Suppose the claim is true for $G_{i}$, and we consider the change of frames equation modulo $R((u))/\mathcal{J}_{i+1}$. We obtain
[…]
@@ 3.4.2. Explicit computations for the Hirzebruch surface $\mathbb{F}_{2}$
Associated to $\mathbb{F}_{2}$ is the fan $\Sigma$ in $N_{\mathbb{R}}=\mathbb{R}^{2}$ whose $1$-dimensional cones are generated by $e_{1}=(1,0)$, $e_{2}=(0,1)$, $e_{3}=(-1,2)$ and $e_{4}=(0,-1)$. Let $D_{i}\subset\mathbb{F}_{2}$ be the toric divisor corresponding to the ray spanned by $e_{i}$. The monoid $P$ is isomorphic to $\mathbb{N}^{2}$ with generators $C_{1}$ and $C_{2}$ corresponding to rational curves lying in $D_{1}$ and $D_{2}$ respectively. Here $C_{1}$ has Chern number $2$ and $C_{2}$ has Chern number 0. The universal piecewise linear function $\varphi:|\Sigma|\rightarrow\mathbb{R}^{2}$ is defined by
@@
Letting $q_{1}=z^{(1,0)}$ and $q_{2}=z^{(0,1)}$, we see that $q_{1}$ has weight degree $2$ and $q_{2}$ has weight degree 0. $X^{\dagger}$ is given as a subvariety of $\text{Spec}(\mathbb{C}[q_{1},q_{2}][z_{1},z_{2},z_{3},z_{4}])$ by the relations
@@
To compare with known results, we will only compute the canonical coordinates and the primitive form when restricted to $\hat{R}=\mathbb{C}[[q_{1},q_{2}]]$ (the small quantum variables), ignoring the other two parameters $t_{1},t_{2}$. The Landau-Ginzburg superpotential is given by $F=z_{1}+z_{2}+z_{3}+z_{4}$. The sheaf $\Omega_{\hat{X}^{{\dagger}}/\hat{S}^{{\dagger}}}^{*}$ of log differential forms is locally free with generators $d\log z_{1}$, $d\log z_{2}$, and the holomorphic volume form is taken to be $\Omega=d\log z_{1}\wedge d\log z_{2}$. The cohomology $H^{*}(\Theta^{*}_{\hat{X}^{{\dagger}}/\hat{S}^{{\dagger}}}[[u]],u\partial+\{F,\cdot\})$ is concentrated at degree 0 which is generated by polynomials in $z_{i}$’s, $q_{i}$’s and $u$.
Consider the action of $u\partial+\{F,\cdot\}$, it leads to the following identities
$(u\partial+\{F,\cdot\})(\tilde{\theta}_{1})$ $=z_{1}-z_{3}=0,$
$(u\partial+\{F,\cdot\})(\tilde{\theta}_{2})$ $=z_{2}+2z_{3}-z_{4}=0,$
@@
They imply the following useful relations:
where $f(q_{2})=\frac{1}{1-4q_{2}}$. We arrange the maximal cones $\sigma_{i}$’s in $\Sigma$ in the counter-clockwise ordering so that $\sigma_{1}=\mathbb{R}_{\geq 0}e_{1}+\mathbb{R}_{\geq 0}e_{2}$. This gives us frames $1$, $z_{3}$, $z_{4}$ and $z_{1}z_{4}$ in $H_{F,+}$ which restrict to $\varphi_{i}$’s in $H_{f,+}$ after modulo $q_{1},q_{2}$, according to the choice made after equation (7).
@@
Next we solve for the sections $DE(\varphi_{i})$’s. We first restrict ourselves to $q_{1}=0$ and then extend to sections in $q_{1}$. Let $\psi_{i}$’s be the restrictions of $DE(\varphi_{i})$’s to $q_{1}=0$. Then we can simply take $\psi_{3}=z_{4}$ and $\psi_{4}=z_{1}z_{4}$. It remains to solve for $\psi_{1}$ and $\psi_{2}$. To do so, we let $\psi_{2}=a(q_{2})z_{3}+b(q_{2})z_{4}+c(q_{2})z_{1}z_{4}$. Then the equation $\nabla_{\frac{\partial}{\partial\log q_{2}}}\psi_{2}=0$ gives us
@@
with $\delta_{i}=\delta_{i}(q_{1},q_{2},u)=\sum_{j=-\infty}^{\infty}\delta_{ij}(q_{1},q_{2})u^{j}$ such that $\delta_{ij}(q_{1},q_{2})\in\mathbb{C}[[q_{1},q_{2}]]$ and the summation starts from $j=-N_{k}$ if we modulo $(q_{1},q_{2})^{k+1}$. The weight degrees of $DE(\varphi_{1})$, $DE(\varphi_{2})$, $DE(\varphi_{3})$ and $DE(\varphi_{4})$ are 0, 22, 22 and $4$ respectively. From these, we know that $\delta_{1}$ has weight degree 0, $\delta_{2}$ and $\delta_{3}$ both have weight degree 2-2 and $\delta_{4}$ has weight degree 4-4. Since the weight degree of $q_{1}$ is $4$ and that of $q_{2}$ is 0, we see that $\delta_{i,j\geq 0}=0$ for $i>0$, $\delta_{0,j>0}=0$ and $\delta_{0,0}$ is independent of $q_{1}$. Modulo $q_{1}$ we find that $\delta_{0,0}\cong 1$ from the expression of $\psi_{1}$. All in all, we see that the primitive form is simply given by $\zeta=1$.